# Statistical Inference for Pairwise Comparison Models

### v2

@@ Abstract
Pairwise comparison models have been widely used for utility evaluation and ranking across various fields. The increasing scale of problems today underscores the need to understand statistical inference in these models when the number of subjects diverges, a topic currently lacking in the literature except in a few special instances. To partially address this gap, this paper establishes a near-optimal asymptotic normality result for the maximum likelihood estimator in a broad class of pairwise comparison models, as well as a non-asymptotic convergence rate for each individual subject under comparison. The key idea lies in identifying the Fisher information matrix as a weighted graph Laplacian, which can be studied via a meticulous spectral analysis. Our findings provide a unified theory for performing statistical inference in a wide range of pairwise comparison models beyond the Bradley–Terry model, benefiting practitioners with theoretical guarantees for their use. Simulations utilizing synthetic data are conducted to validate the asymptotic normality result, followed by a hypothesis test using a tennis competition dataset.
Keywords: Imbalanced data, Maximum likelihood estimation, Ordinal data, Pairwise comparison models, Sparsity
@@ 1 Introduction
Pairwise comparison involves assessing subjects in pairs to establish their relative preferences, a practice pertinent to numerous applications such as sports analytics (Glickman and Jones, 1999; Ley et al., 2019; Tsokos et al., 2019), econometrics (McFadden, 1973), video coding (Lee et al., 2011), and social science (Loewen et al., 2012). A prevalent approach for pairwise comparison modeling employs a latent score framework. Originated from the ideas in Thurstone (Thurstone, 1927) and Zermelo (Zermelo, 1929), a mathematical model for pairwise comparison data analysis was proposed by Bradley and Terry (Bradley and Terry, 1952). Since then, multiple generalizations have been developed, including ordinal models such as the Rao–Kupper model (Rao and Kupper, 1967) and the Davidson model (Davidson, 1970), which account for ties, the cumulative link model (Agresti, 1992) that considers more refined ordinal scales, and cardinal models such as the paired cardinal model (Shah et al., 2016). We recommend Cattelan (2012) for a review of pairwise comparison modeling from a practical perspective.
@@
In light of the fast-growing scale of problems in the big-data era, a recent research area of pairwise comparison models focuses on understanding the asymptotics of estimating the latent score vector as the number of compared subjects approaches infinity. Along this line of work, considerable attention has been devoted to the study of the Bradley–Terry (BT) model, which is popular due to its simple parametrization form and connections to other problems such as matrix balancing (Qu et al., 2023) and completion (Chen et al., 2023). Notably, an efficient Minorize-Maximization (MM) algorithm for implementing the maximum likelihood estimator (MLE) was proposed in Hunter (2004) and accelerated in Newman (2023) in the BT model, and its uniform consistency was established when the comparison graph is dense (Simons and Yao, 1999; Yan et al., 2012). Subsequent works have established analogous consistency results for alternative estimators such as the spectral estimator and regularized MLE in the sparse settings under various error metrics (Chen and Suh, 2015; Shah et al., 2016; Negahban et al., 2017; Chen et al., 2019). The first result that generalized the uniform consistency of the MLE to the sparse setting appeared in Han et al. (2020), paralleled by a few other works with a slightly different focus (e.g. asymptotic efficiency and minimax rates); see Hendrickx et al. (2020); Bong and Rinaldo (2022); Chen et al. (2022). It is noteworthy that Han et al. (2020) also contains an asymptotic normality result for dense graphs, which was further explored in the sparse setting (Liu et al., 2023; Gao et al., 2023).
@@
Despite extensive research on the asymptotic properties of the BT model, a significant gap persists between theory and practice, as many comparison models used in applications either have multiple scales or a continuous outcome. This suggests a practical need to extend this understanding to other widely used yet less explored models. The recent work Han et al. (2023) partially addresses this by introducing a general pairwise comparison model framework and proving a uniform consistency result for the MLE therein. However, an asymptotic normality result has not yet been established. Nonetheless, such a result is often crucial for practical statistical inference and cannot be deduced from the existing approaches to analyzing the BT model.
@@
  • We verify that the MLE in many pairwise comparison models in the literature, including the BT model, the Thurstone–Mosteller model (Thurstone, 1927; Mosteller, 1951), the Davidson model, the Rao–Kupper model, the cumulative link model, and the paired cardinal model, is asymptotically normal under near-optimal sparsity conditions. We also provide explicit and computable formulas for their asymptotic variance that may be of interest to practitioners using these models.
@@
Simons and Yao (1999) BT homogeneous $\Omega(1)$
Yan et al. (2012) BT heterogeneous $\Omega(1)$
Han et al. (2020) BT homogeneous $\widetilde{\Omega}(n^{-1/10})$
Gao et al. (2023) BT homogeneous $\widetilde{\Omega}(n^{-1})$
@@
Table 1: Comparison of several asymptotic normality results in the pairwise comparison model literature. Homogeneous comparison graphs indicate the balanced data while heterogeneous graphs correspond to the imbalanced data. The notation $\Omega(\cdot)$ is the standard Bachmann–Landau notation and $\widetilde{\Omega}(\cdot)$ means that the asymptotic relation holds up to polynomial terms of $\log n$.
The rest of the paper is organized as follows. In Section 2, we introduce the mathematical framework for general pairwise comparison models. In Section 3, we establish the main results concerning asymptotic normality and non-asymptotic convergence rates (Theorems 8, 11, 12) for the MLE and provide the relevant interpretations. In Section 4, we show that the MLE in many pairwise comparison models in the literature is asymptotically normal under minimal sparsity assumptions. Furthermore, we offer explicit formulas for their asymptotic variance. In Section 5, we validate the theoretical findings using both synthetic and real data. In Section 6, we outline the key ideas for the proof of the main result, deferring the technical details to the appendix section. In Section 7, we point out some future directions.
@@ 2 General pairwise comparison models
We introduce the general pairwise comparison models proposed in Han et al. (2023). Consider an undirected graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ with $n$ vertices, where ${\mathcal{V}}=[n]:=\{1,\ldots,n\}$ represents the subjects being modeled, and ${\mathcal{E}}$ is the edge set of the comparison data, that is, for any $i,j\in{\mathcal{V}}$, $(i,j)\in{\mathcal{E}}$ if there exists a comparison between subjects $i$ and $j$. For ease of illustration, we assume that ${\mathcal{G}}$ is simple so that there exists at most one edge between two subjects; the general case consisting of multiple edges can be considered similarly but requires heavier notation. We use $\delta\{i\}=\{j\in{\mathcal{V}}:(i,j)\in{\mathcal{E}}\}$ to denote the graph neighbourhood of $i$, which refers to the set of vertices adjacent to $i$. For the moment, we assume that ${\mathcal{G}}$ is known and describe the statistical model for the comparison outcomes.
@@
Definition 1 (Valid parameterization)
@@
(A2). (symmetry) $f$ is even with respect to $(x;y)$, that is, $f(x;y)=f(-x;-y)$, for $(x,y)\in{\mathbb{A}}\times{\mathbb{R}}.$
@@
Having specified $f$, we can write down the likelihood function given the observed data and compute the MLE $\widehat{{\bm{u}}}$ for ${\bm{u}}^{*}$ under identifiability constraints; see Section 3. It was shown in Han et al. (2023) that $\widehat{{\bm{u}}}$ is a uniformly consistent estimator for ${\bm{u}}^{*}$ under appropriate conditions on both $f$ and ${\mathcal{G}}$. To identify graph configurations satisfying these conditions, it is often convenient to assume that ${\mathcal{G}}$ is sampled from certain random graph ensembles. A commonly used one is the random graph model ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$ that generalizes the Erdős–Rényi model (Erdős and Rényi, 1960).
Definition 2 (Random graph models)
@@
In the following section, we assume that the comparison graph ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ is sampled from ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$. Meanwhile, for every $(i,j)\in{\mathcal{E}}$, we assume that a single comparison outcome $X_{ij}\sim f(x;u^{*}_{i}-u^{*}_{j})$ is observed. Note that $X_{ij}=-X_{ji}$, and we assume that $\{X_{ij}\}_{i<j,(i,j)\in{\mathcal{E}}}$ are independent. It is worth noting that we allow $\lim\sup_{n\to\infty}q_{n}/p_{n}=+\infty$, which leads to heterogeneity of the observed data. In particular, one subject may have $nq_{n}$ comparisons while another has only $np_{n}$ comparisons. Such an imbalance of data volume will bring additional difficulty in theoretical data analysis.
@@ 3 Main results
Note that $l({\bm{u}})$ is invariant if all components of ${\bm{u}}$ are translated by a constant. To ensure model identifiability, we let ${\bm{1}}_{n}^{\top}{\bm{u}}=0$, where ${\bm{1}}_{n}$ is the all-ones vector with the same dimension as ${\bm{u}}$. The MLE ${\widehat{\bm{u}}}$ for ${\bm{u}}^{*}$ satisfies
${\widehat{\bm{u}}}=\operatorname*{arg\,max}_{{\bm{u}}\in{\mathbb{R}}^{n}:{\bm{1}}_{n}^{\top}{\bm{u}}=0}l({\bm{u}}).$ (1)
To state the asymptotic normality result for $\widehat{{\bm{u}}}$, we need the following assumptions. Let $M_{n}=\max_{i,j\in[n]}|u^{*}_{i}-u^{*}_{j}|$ denote the dynamic range of ${\bm{u}}^{*}$. The global discrepancy of the model, which measures the probability of a subject with the highest score winning over another with the lowest score, is defined as
@@
$\frac{\log n}{np_{n}|\log c_{n,1}|}\to 0\qquad n\to\infty.$ (2)
@@
where $\|\cdot\|_{\psi_{2}}$ is the subgaussian norm (Vershynin, 2018, Definition 2.5.6).
@@
Assumptions 4-6 are sufficient to guarantee the uniform consistency of the maximum likelihood estimator (Han et al., 2023). The common pathway from consistency to asymptotic normality is via Taylor’s expansion. As a result, one often require more stringent regularity conditions on higher-order derivatives.
Assumption 7
@@
Theorem 8 (Asymptotic normality)
@@
$\beta_{n}=\max\left\{\frac{c_{n,2}^{2}c_{n,4}^{5/2}c_{n,5}}{c_{n,3}^{5}},\frac{c_{n,2}c_{n,4}^{11/2}}{c_{n,3}^{6}}\right\}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty,$ (3)
then for each $i\in[n]$, $\{\rho_{i}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i}-u_{i}^{*})\to N(0,1)$, where $N(0,1)$ is the standard normal distribution and the asymptotic variance $\rho_{i}({\bm{u}}^{*})$ is given by
$\rho_{i}({\bm{u}}^{*})$ $=\left[\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\frac{\{\partial_{2}f(x;u^{*}_{i}-u^{*}_{j})\}^{2}}{f(x;u^{*}_{i}-u^{*}_{j})}\operatorname{d}\!{x}\right]^{-1}.$ (4)
@@
Remark 9
Subject to additional conditions, such as a uniform upper bound on the number of edges between two subjects that is independent of ${\bm{u}}^{*}$, Theorem 8 still holds when two subjects are compared multiple times. Under such circumstances, the asymptotic variance formula (4) remains valid if one counts multiple edges for each adjacent vertex.
@@
Corollary 10 can be further simplified under balanced data conditions. If $\sup_{n}q_{n}/p_{n}<\infty$, then the condition in Corollary 10 reduces to $(\log n)^{8}/(np_{n})\to 0$ as $n\to\infty$. This bound coincides with the well-known Erdős–Rényi connectivity threshold up to logarithmic factors and is thus close to optimal. In general, Corollary 10 allows $p_{n}$ and $q_{n}$ to have different asymptotic orders. For instance, if $p_{n}=n^{-b}$ for $10\leq b<$ 1, then choosing $q_{n}=n^{-b^{\prime}}$ for any $b^{\prime}\geq 0$ with $b-(1-b)/10<b^{\prime}\leq b$ satisfies the required conditions. To the best of our knowledge, this is the first result that works for general comparison models with imbalanced data.
Theorem 8 provides the limit distribution of the MLE, which lays the foundation for further statistical inference, such as the construction of interval-based estimates, testing hypothesis on the difference of merits, and testing the hypothesis on ranking position (Fan et al., 2024). Additionally, we can derive a non-asymptotic error rate for each parameter as well as a practical construction of interval-valued estimators, which are summarized in Theorem 11 and Theorem 12, respectively.
Theorem 11 (Individual estimation error)
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. If both Assumptions 47 and (3) hold, then there exists an absolute constant $C>0$ such that for each $i\in[n]$, the following holds for all sufficiently large $n$ with probability tending to one:
$|\widehat{u}_{i}-u_{i}^{*}|\leq C\cdot\frac{c_{n,2}}{c_{n,3}}\left(\frac{\log n}{|\delta\{i\}|}\right)^{1/2}.$ (5)
We call the result in (5) the individual estimation error bound. Compared with the existing works (Han et al., 2020; Chen et al., 2022), this result presents new findings on the individual estimation error. Specifically, previous works considered the overall estimation error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$, while our result in (5) further illustrates the error for each individual. As expected, this error depends on the degree $|\delta\{i\}|$, which serves as a local normalization factor for each individual. When the comparison graph is homogeneous, individual estimation errors converge at the same rate, matching the results in Han et al. (2020); Chen et al. (2022). For heterogeneous graphs, however, individuals will have different convergence rates, which cannot be reflected in the worst-case error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$.
@@
$\rho_{i}({\widehat{\bm{u}}})$ $=\left[\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\frac{\{\partial_{2}f(x;\widehat{u}_{i}-\widehat{u}_{i})\}^{2}}{f(x;\widehat{u}_{i}-\widehat{u}_{i})}\operatorname{d}\!{x}\right]^{-1}.$ (6)
@@
Theorem 12 (Confidence interval)
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 47, if (3) holds, then for any fixed $i\in[n]$ and $\alpha\in(0,1)$,
@@
The computation of $\mathcal{CI}_{i}$ is straightforward after obtaining ${\widehat{\bm{u}}}$. For example, if one wants to construct a $95\%$-confidence interval for the $i$th subject, one can take $\mathcal{CI}_{i}=\big{(}\widehat{u}_{i}-1.96\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2},\widehat{u}_{i}+1.96\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2}\big{)}$.

### v3

@@ Abstract
Pairwise comparison models have been widely used for utility evaluation and rank aggregation across various fields. The increasing scale of modern problems underscores the need to understand statistical inference in these models when the number of subjects diverges, a topic that is currently underexplored in the literature. To address this gap, this paper establishes a near-optimal asymptotic normality result for the maximum likelihood estimator in a broad class of pairwise comparison models. The key idea lies in identifying the Fisher information matrix as a weighted graph Laplacian, which can be studied via a meticulous spectral analysis. Our findings provide theoretical foundations for performing statistical inference in a wide range of pairwise comparison models beyond the Bradley–Terry model. Simulations utilizing synthetic data are conducted to validate the asymptotic normality result, followed by a hypothesis test using a tennis competition dataset.
Keywords: Confidence interval; Maximum likelihood estimation; Ordinal data; Pairwise comparison; Sparsity.
@@ 1 Introduction
Pairwise comparison involves assessing subjects in pairs to establish their relative preferences, a practice pertinent to various applications such as econometrics (eco1958; mcfadden1973conditional; 10.1257/pandp.20221063; lentz2023anatomy), sports analytics (baker2014dynamic; bozoki2016application; angelini2022weighted; collingwood2022evaluating), social science (loewen2012testing; 10.1111/rssa.12124; wapman2022quantifying), and, more recently, human preference studies aimed at improving models in artificial intelligence (Christiano2017; Rafailov2023; Sun2025). A prevalent approach for pairwise comparison modeling employs a latent score framework. Originated from the ideas in Thurstone (thurstone1927method) and Zermelo (MR1545015), a mathematical model for pairwise comparison data analysis was formulated by Bradley and Terry (MR0070925). Since then, multiple generalizations have been developed, including ordinal models such as the Rao–Kupper model (MR217963) and the Davidson model (davidson1970extending), which account for ties, the cumulative link model (agresti1992analysis) that considers more refined ordinal scales, and cardinal models such as the paired cardinal model (MR3504618). We recommend MR3012434 for a review of pairwise comparison modeling from a practical perspective.
@@
In light of the fast-growing scale of problems in the big-data era, a recent research area of pairwise comparison models focuses on understanding the asymptotics of estimating the latent score vector as the number of compared subjects approaches infinity. Along this line of work, considerable attention has been devoted to the study of the Bradley–Terry (BT) model, which is popular due to its simple parametrization form and connections to other problems such as matrix balancing (qu2023sinkhorn) and completion (chen2023note). Notably, an efficient Minorize-Maximization (MM) algorithm for implementing the maximum likelihood estimator (MLE) was proposed in MR2051012 and accelerated in newman2023efficient in the BT model, and its uniform consistency was established when the comparison graph is dense (MR1724040; MR2987494). Subsequent works have established analogous consistency results for alternative estimators such as the spectral estimator and regularized MLE in the sparse settings under various error metrics (chen2015spectral; MR3504618; MR3613103; MR3953449). Recently, han2020asymptotic and chen2022optimal extended the uniform consistency of the MLE to the sparse setting, paralleled by a few other works with a slightly different focus (e.g., asymptotic efficiency and minimax rates); see hendrickx2020minimax; bong2022generalized. It is noteworthy that han2020asymptotic also contains an asymptotic normality result for dense graphs, which was further explored in the sparse setting (MR4560195; gao2023uncertainty).
@@
Despite extensive research on the asymptotic properties of the BT model, a gap persists between theory and practice as many comparison models used in applications either have multiple scales or a continuous outcome. This suggests a practical need to extend this understanding to other widely used yet less explored models. The recent work (han2022general) partially addresses this by introducing a general pairwise comparison model framework and proving a uniform consistency result for the MLE therein. However, an asymptotic normality result has not yet been established. Such a result is often crucial for practical inference and cannot be deduced from the existing approaches to analyzing the BT model.
@@
  • We verify that the MLE in many pairwise comparison models in the literature, including the BT model, the Thurstone–Mosteller model (thurstone1927method; mosteller2006remarks), the Davidson model, the Rao–Kupper model, the cumulative link model, and the paired cardinal model, is asymptotically normal under near-optimal sparsity conditions. We also provide explicit and computable formulas for their asymptotic variance that may be of interest to practitioners using these models.
@@
MR1724040 BT homogeneous $\Omega(1)$
MR2987494 BT heterogeneous $\Omega(1)$
han2020asymptotic BT homogeneous $\widetilde{\Omega}(n^{-1/10})$
gao2023uncertainty BT homogeneous $\widetilde{\Omega}(n^{-1})$
@@
[…]
The rest of the paper is organized as follows. In Section 2, we introduce the mathematical framework for general pairwise comparison models. In Section 3, we establish the main results concerning asymptotic normality and non-asymptotic convergence rates (Theorems 1, 2, 3) for the MLE and provide the relevant interpretations. In Section 4, we show that the MLE in many pairwise comparison models in the literature is asymptotically normal under minimal sparsity assumptions. Furthermore, we offer explicit formulas for their asymptotic variance. In Section 5, we validate the theoretical findings using both synthetic and real data. In Section 6, we outline the key ideas for the proof of the main result, deferring the technical details to the appendix section. In Section 7, we point out some future directions.
@@ 2 General pairwise comparison models
We introduce the general pairwise comparison models proposed in han2022general. Consider an undirected graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ with $n$ vertices, where ${\mathcal{V}}=[n]:=\{1,\ldots,n\}$ represents the subjects being modeled, and $\mathcal{E}$} is the edge set of the comparison data, that is, for any $i,j\in{\mathcal{V}}$, $(i,j)\in{\mathcal{E}}$ if there exists a comparison between subjects $i$ and $j$. For ease of illustration, we assume that $\mathcal{G}$} is simple so that there exists at most one edge between two subjects; the general case consisting of multiple edges can be considered similarly but requires heavier notation. We use $\delta\{i\}=\{j\in{\mathcal{V}}:(i,j)\in{\mathcal{E}}\}$ to denote the graph neighbourhood of $i$, which refers to the set of vertices adjacent to $i$. For the moment, we assume that $\mathcal{G}$} is known and describe the statistical model for the comparison outcomes.
@@
Definition 1 (Valid parameterization).
@@
(A2). (symmetry) $f$ is even with respect to $(x;y)$, that is, f(x;y)=f(x;y)f(x;y)=f(-x;-y), for $(x,y)\in{\mathbb{A}}\times{\mathbb{R}}.$
@@
Having specified $f$, we can write down the likelihood function given the observed data and compute the MLE $\widehat{{\bm{u}}}$ for $\bm{u}$$^{*}$ under identifiability constraints; see Section 3. It was shown in han2022general that $\widehat{{\bm{u}}}$ is a uniformly consistent estimator for $\bm{u}$$^{*}$ under appropriate conditions on both $f$ and $\mathcal{G}$}. To identify graph configurations satisfying these conditions, it is often convenient to assume that $\mathcal{G}$} is sampled from certain random graph ensembles. A commonly used one is the random graph model ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$ that generalizes the Erdős–Rényi model (MR0125031).
Definition 2 (Random graph models).
@@
In the following section, we assume the comparison graph ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ is sampled from ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$. Meanwhile, for every $(i,j)\in{\mathcal{E}}$, we assume that a single comparison outcome $X_{ij}\sim f(x;u^{*}_{i}-u^{*}_{j})$ is observed. Note that $X_{ij}=-X_{ji}$, and we assume that $\{X_{ij}\}_{i<j,(i,j)\in{\mathcal{E}}}$ are independent. It is worth noting that we allow $\lim\sup_{n\to\infty}q_{n}/p_{n}=+\infty$, which leads to heterogeneity of the observed data. In particular, one subject may have $nq_{n}$ comparisons while another has only $np_{n}$ comparisons. Such an imbalance of data volume will bring additional difficulty in theoretical data analysis.
@@ 3 Main results
Note that $l({\bm{u}})$ is invariant if all components of $\bm{u}$} are shifted by a constant. To ensure model identifiability, we let ${\bm{1}}_{n}^{\top}{\bm{u}}=0$, where ${\bm{1}}_{n}$ is the all-ones vector with the same dimension as $\bm{u}$}. The MLE $^{\widehat{\bm{u}}}$ for $\bm{u}$$^{*}$ satisfies
${\widehat{\bm{u}}}=\operatorname*{arg\,max}_{{\bm{u}}\in{\mathbb{R}}^{n}:{\bm{1}}_{n}^{\top}{\bm{u}}=0}l({\bm{u}}).$ (3.1)
To state the asymptotic normality result for $\widehat{{\bm{u}}}$, we need the following assumptions. Let
$M_{n}=\max_{i,j\in[n]}|u^{*}_{i}-u^{*}_{j}|$
denote the dynamic range of $\bm{u}$$^{*}$. The global discrepancy of the model, which measures the probability of a subject with the highest score winning over another with the lowest score, is defined as
@@
$\frac{\log n}{np_{n}|\log c_{n,1}|}\to 0\qquad n\to\infty.$ (3.2)
@@
where $\|\cdot\|_{\psi_{2}}$ is the subgaussian norm (vershynin2018high, Definition 2.5.6).
@@
Assumptions 1-3 are sufficient to guarantee the uniform consistency of the maximum likelihood estimator (han2022general). The common pathway from consistency to asymptotic normality is via Taylor’s expansion. As a result, one often require more stringent regularity conditions on higher-order derivatives.
Assumption 4.
@@
Theorem 1 (Asymptotic normality).
@@
$\beta_{n}=\max\left\{\frac{c_{n,2}^{2}c_{n,4}^{5/2}c_{n,5}}{c_{n,3}^{5}},\frac{c_{n,2}c_{n,4}^{11/2}}{c_{n,3}^{6}}\right\}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty,$ (3.3)
then for each $i\in[n]$, $\{\rho_{i}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i}-u_{i}^{*})\to\mathcal{N}(0,1)$, where $\mathcal{N}(0,1)$ is the standard normal distribution and the asymptotic variance $\rho_{i}({\bm{u}}^{*})$ is given by
$\rho_{i}({\bm{u}}^{*})$ $=\left[\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\frac{\{\partial_{2}f(x;u^{*}_{i}-u^{*}_{j})\}^{2}}{f(x;u^{*}_{i}-u^{*}_{j})}\operatorname{d}\!{x}\right]^{-1}.$ (3.4)
@@
Remark 2.
Subject to additional conditions, such as a uniform upper bound on the number of edges between two subjects that is independent of $\bm{u}$$^{*}$, Theorem 1 still holds when two subjects are compared multiple times. Under such circumstances, the asymptotic variance formula (3.4) remains valid if one counts multiple edges for each adjacent vertex.
@@
Corollary 1 can be further simplified under balanced data conditions. If $\sup_{n}q_{n}/p_{n}<\infty$, then the condition in Corollary 1 reduces to $(\log n)^{8}/(np_{n})\to 0$ as $n\to\infty$. This bound coincides with the well-known Erdős–Rényi connectivity threshold up to logarithmic factors and is thus close to optimal. In general, Corollary 1 allows $p_{n}$ and $q_{n}$ to have different asymptotic orders. For instance, if $p_{n}=n^{-b}$ for $0\leq b<1$, then choosing $q_{n}=n^{-b^{\prime}}$ for any $b^{\prime}\geq 0$ with $b-(1-b)/10<b^{\prime}\leq b$ satisfies the required conditions. To the best of our knowledge, this is the first result that works for general comparison models with imbalanced data.
Theorem 1 provides the limiting distribution of the MLE, which lays the foundation for further statistical inference, such as the construction of interval-based estimates, testing hypothesis on the difference of merits, and testing the hypothesis on ranking position (doi:10.1080/01621459.2024.2316364). Additionally, we can derive a non-asymptotic error rate for each parameter as well as a practical construction of interval-valued estimators, which are summarized in Theorem 2 and Theorem 3, respectively.
Theorem 2 (Individual estimation error).
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. If both Assumptions 14 and (3.3) hold, then there exists an absolute constant $C>0$ such that for each $i\in[n]$, the following holds for all sufficiently large $n$ with probability tending to one:
$|\widehat{u}_{i}-u_{i}^{*}|\leq C\cdot\frac{c_{n,2}}{c_{n,3}}\left(\frac{\log n}{|\delta\{i\}|}\right)^{1/2}.$ (3.5)
We call the result in (3.5) the individual estimation error bound. Compared with the existing works (han2020asymptotic; chen2022optimal), this result presents new findings on the individual estimation error. Specifically, previous works considered the overall estimation error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$, while our result in (3.5) further illustrates the error for each individual. As expected, this error depends on the degree $|\delta\{i\}|$, which serves as a local normalization factor for each individual. When the comparison graph is homogeneous, individual estimation errors converge at the same rate, matching the results in han2020asymptotic and chen2022optimal. For heterogeneous graphs, however, individuals will have different convergence rates, which cannot be reflected in the worst-case error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$.
@@
[…]
@@
Theorem 3 (Confidence interval).
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 14, if (3.3) holds, then for any fixed $i\in[n]$ and $\alpha\in(0,1)$,
@@
The computation of $\mathcal{CI}_{i}$ is straightforward after obtaining $^{\widehat{\bm{u}}}$. For example, if one wants to construct a $95\%$-confidence interval for the $i$th subject, one can take $\mathcal{CI}_{i}=\big(\widehat{u}_{i}-1.96\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2},\widehat{u}_{i}+1.96\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2}\big)$.